OpenStudy (anonymous):

Need to factor t^2+4z^2

OpenStudy (anonymous):

do you know complexx numbers ?!

OpenStudy (anonymous):

No I don't

OpenStudy (anonymous):

If this what you mean \[t^2+4z^2\] it can't be factorized

OpenStudy (anonymous):

Please let me know if this is correct: (t+2z)(t+2z) factor completed

OpenStudy (anonymous):

in other words it would be written prime... correct?

OpenStudy (anonymous):

Sorry , but I didn't understaand what you really want to do ! :)

OpenStudy (anonymous):

ok I have another one for you. I am asked to factor this polynomial: y^3+y^2-4y-4

OpenStudy (anonymous):

okk ! this ! it can be factorized :) ! you may know that\[(-1)^3+(-1)^2-4(-1)-4=0\] so , this "polynomial" can be factorized by (y--1)=(y+1)

OpenStudy (anonymous):

what would be considered the common factor group?

OpenStudy (anonymous):

kk you didn't undersand what I said ! It's normal ! we will try to do this in other way ! kk ?!

OpenStudy (anonymous):

This is what I has before: y^3+y^2-4y-4 y(y^2+y) -4(y+1) (y^2+y)-(y-4) (y^2+y)(y-4)(4+-1)

OpenStudy (anonymous):

ok. trust me it has been a reeeallllyyy long time since I have done algebra and my brain is hurting over this lol

OpenStudy (anonymous):

I have to be able to show my steps as well as explain it. At this point I am just trying to figure out how to do this :)

OpenStudy (anonymous):

no no ! good ! you were in the right way ! \[y^3+y^2=y^2(y+1)\] -4y-4=-4(y+1)

OpenStudy (anonymous):

you're reallyy a smart person !

OpenStudy (anonymous):

I seriously doubt that right now lol

OpenStudy (anonymous):

no no ! you're doing good trust me !

OpenStudy (anonymous):

How would I go about explaining how I arrived at this conclusion?

OpenStudy (anonymous):

so would my common factor be the (y+1)?

OpenStudy (anonymous):

yeaaaaah !

OpenStudy (anonymous):

now give me the final answer !

OpenStudy (anonymous):

(y+1)(y+1)(y+-4)

OpenStudy (anonymous):

I have to utilize grouping because of the four terms right

OpenStudy (anonymous):

give me another answer :'( !

OpenStudy (anonymous):

\[y^3+y^2-4y+4=y^2(y+1)-4(y+1)\] now (y+1) is the common factor ! kk

OpenStudy (anonymous):

go on what are waiting for ?! :) :)

OpenStudy (anonymous):

differences in squares would be: (y+1)(y^2+4)?

OpenStudy (anonymous):

(y+1)(y+2)(y+-2)?

OpenStudy (anonymous):

completely factored at (y+1)(y^2+2)(y+-2)

OpenStudy (anonymous):

now you're good ! But you wrote (y^2+4) correct this the final ansswer is correct !

OpenStudy (anonymous):

so is (y+1)(y^2+2)(y^-2) correct?

OpenStudy (anonymous):

yeaaaah that's correct !

OpenStudy (anonymous):

thank you !!!! :)

OpenStudy (anonymous):

You're welcome !anytime !

OpenStudy (anonymous):

o wait shouldn't it be like this: (y+1)(y+2)(y-2)

OpenStudy (anonymous):

final factorization that is?

OpenStudy (anonymous):

(y+1)(y+2)(y+-2) (Ugh)

OpenStudy (anonymous):

okk I didn't see well ! it should be ! the right answer is (y+1)(y+2)(y-2) which is the same as (y+1)(y+2)(y+-2)

OpenStudy (anonymous):

oh ok lol

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

yw ! :D !

OpenStudy (anonymous):

and good luck !

OpenStudy (anonymous):

thank you because boy am I going to need it

OpenStudy (anonymous):

:) yeaaah we all need luck !